IMPROVED PARAMETERIZATION FOR DIATOMIC BORN-OPPENHEIMER BREAKDOWN EFFECTS, AND A NEW COMBINED-ISOTOPES ANALYSIS FOR HD AND DF

\underline{ROBERT J. LE ROY}, Guelph-Waterloo Centre for Graduate Work in Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

Although the Ross et al.^{a}-Bunker^{b}-Watson^{c} expression for the vibration-rotation levels of isotopomers-α of diatomic molecule A–B, E^{α}(v, J) = ∑_{l,m} \frac{U_{l,m}}{(μ_{α})^{m+l/2}} (1 + \frac{m_{e}}{m_{A}^{α}} Δ_{l,m}^{A} + \frac{m_{e}}{m_{B}^{α}} Δ_{l,m}^{B}) (v + 1/2)^{l} [J(J + 1)]^{m} , is fairly widely used, it has a number of deficiencies. In particular: (i) it is not clear how to treat the (l, m) = (0, 0) term, (ii) the U_{l,m} constants have inconveniently varying units depending on fractional powers of the mass, (iii) fits to data for only two isotomers cannot yield “true” mass-independent U_{l,m} values, (iv) for non-hydrides, the magnitudes of the U_{l,m} constants are not even qualitatively similar to those of the analogous familiar Dunham Y_{l,m} constants, and (v) direct fits to this expression are impossible, so non-linear fits or indirect methods must be used to determine the Δ_{l,m}^{a} values and their uncertainties. Moreover, this formulation is incompatible with the use of non-Dunham expressions (such as neardissociation expansions) for the vibrational energies and rotational and centrifugal distortion constants. The present paper presents an alternate description of the Born-Oppenheimer and JWKB breakdown corrections which resolves all of these difficulties, and illustrates its use by a combined-isotopes analysis of the best available microwave, infrared and electronic data for the ground electronic states of HF and DF.

aA.H.M. Ross, R.S. Eng and H. Kildal, Opt. Commun. 12, 433 (1974)

bP.R. Bunker, J. Mol. Spectrosc. 68, 367 (1977)

cJ.K.G. Watson J. Mol. Spectrosc. 68, 367 (1977)